OpenStudy (anonymous):

1. A gardener is planning to plant a rectangular vegetable garden during the spring. The rectangle below represents a plan for the garden with a width of x feet. area of the rectangle is x^2 + 12x length =l width =x

OpenStudy (anonymous):

@4everaddicted2anime

OpenStudy (anonymous):

the question is, b. Find the dimensions (length × width) and the area of the garden if x = 10 feet. (Hint: Substitute 10 for x in the expressions for length, width, and area.)

OpenStudy (anonymous):

i only need b this time.

OpenStudy (4everaddicted2anime):

What did you find for w?

OpenStudy (anonymous):

i have some of the work from a if you want me to post is

OpenStudy (anonymous):

it

OpenStudy (4everaddicted2anime):

Can you? I can't find any solutions to it with using complex numbers

OpenStudy (anonymous):

x * l = x^2+12x x * l = x (x +12) l = x+12 x = 10 feet l = x +12 = 10 + 12 = 22 feet A = x^2+12x = (10)^2 + 12(10) = 100 + 120 = 220 feet^2

OpenStudy (4everaddicted2anime):

That is the area

OpenStudy (4everaddicted2anime):

okay so w is equal to x

OpenStudy (4everaddicted2anime):

Do you understand how I got that?

OpenStudy (anonymous):

no

OpenStudy (4everaddicted2anime):

For the factored version of this equation you wrote x(x+12) (X+12) is the length and whatever is leftover is the width

OpenStudy (anonymous):

ok

OpenStudy (4everaddicted2anime):

There ya go. Part B is answered

OpenStudy (anonymous):

ok so whats the width and the leangth

OpenStudy (anonymous):

22= leangth

OpenStudy (4everaddicted2anime):

yes

OpenStudy (4everaddicted2anime):

and width is the value of x(10)

OpenStudy (anonymous):

100

OpenStudy (4everaddicted2anime):

what I meant was the value of x is 10, so the width is 10

OpenStudy (anonymous):

ok and the area is 220 right

OpenStudy (4everaddicted2anime):

Yes